# Semi-vortex solitons and their excited states in spin-orbit-coupled binary bosonic condensates

### v1

@@ I Introduction
Various methods have been proposed to achieve this purpose. On the one hand, it has been predicted that long-range interactions in BEC of dipole atoms, carrying permanent magnetic moments, can support stable 2D or even 3D matter-wave solitons [30, 31, 32, 33, 34]. Another possibility is to use the combination of the spin-orbit coupling (SOC) between two components of binary BEC and attractive mean-field nonlinearity [35, 36, 37, 38, 39, 40, 41, 42]. However, the stabilization provided by these systems is limited to the fundamental (ground state, GS) solitons. In particular, the SOC system makes it possible to find 2D excited-state (ES) solitons, which, however, are completely unstable [35]. Naturally, ES solitons exhibit a richer amplitude and phase structure than their GS counterparts, which makes stabilization of ES solitons another relevant problem, that has not been resolved in previous works [26, 43].
@@
Our analysis demonstrates that the ES solitons may be stable, at least, up to $S_{+}=6$. Systematically collected numerical results reveal that the ES solitons with the vorticity in both components exist provided that their total does not exceed a collapse threshold $N_{thr}$, which surpasses the well-known threshold ($N_{c}\approx 5.85$) for the fundamental (Townes) single-component solitons in the free 2D space [21, 22]. The value $N_{thr}$ increases with the growth of the additional vorticity $S_{+}$. These findings offer, for the first time, a solution of the problem of the creation of stable ES solitons in SOC systems.
The following presentation is structured as follows. In Section 2, we introduce the model, based on a system of two spin-orbit-coupled Gross-Pitaevskii equations (GPEs), with the opposite signs of the cubic nonlinearity in the two components. In Section 3, we produce SV soliton solutions of the GS and ES types and test their stability by means of numerical methods. In Section 4, we address moving SV solitons and their stability, which is a nontrivial problem, as the SOC breaks the system’s Galilean invariance. A critical velocity $V_{\mathrm{cr}}$, up to which the stable moving solitons exist, is identified as a function of the total norm $N$. Finally, Section 5 concludes the paper.
@@ II The model
where positive $g_{+}$ and $g_{-}$ coefficients are strengths of the self-attraction and repulsion in the components $\Psi_{+}$ and $\Psi_{-}$, respectively, and $\gamma$ is the cross-interaction strength. The linearized form of Eqs. (2) and (3) for plane-wave solutions, $\Psi_{\pm}\sim\exp\left(-i\mu_{\mathrm{linear}}t+ik_{x}x+ik_{y}y\right)$, yields the dispersion relation
@@
Obviously, $N$ is a dynamical invariant of the system of Eqs. (2) and (3). In the analysis following below, we treat $(N,g_{+},g_{-},\gamma)$ as a set of the system’s control parameters.
@@
is also a dynamical invariant. Note that the  substitution of ansatz (6) in expression (13) yields a simple relation between the angular momentum and norm (10) of the stationary states, which applies equally well to the the GS and ESs:
@@ III.1 Ground-state (GS) solitons
For two typical values of the total norm, $N=4.5$ and $N=3.5$, properties of the GS solutions are summarized in Figs. 2(a,b,c), where the chemical potential, effective radius, and the norm ratio of the two components, defined as per Eq. (9), are plotted vs. the cross-interaction strength $\gamma$ in Eqs. (2) and (3). In these plots, $\gamma$ takes both positive and negative values, which correspond, respectively, to the cross-attraction and repuslion, while the self-interaction coefficients are fixed as $g_{+}=g_{-}=1$. In Fig. 2(a), the SV soliton suffers delocalization, turning into an infinitely broad state with an infinitely small amplitude, as chemical potential $\mu$ is reaching the limit value (5), $\mu_{\mathrm{cutoff}}=-1/2$, at $\gamma\rightarrow\gamma_{\mathrm{cr}}\left(N=4.5\right)=-0.775$ or $\gamma_{\mathrm{cr}}\left(N=3.5\right)=-0.24$. For other values of the norm (in the interval of $5.53<N<$ 5.5), the dependence of $\gamma_{\mathrm{cr}}$ on $N$ is plotted in Fig. 2(d). Self-trapped bound states (solitons) do not exist if $\gamma$ is “too negative”, taking values $\gamma<\gamma_{\mathrm{cr}}(N)$, as the self-trapping is obviously impossible when the overall nonlinearity is self-repulsive. The transition to the delocalization is also demonstrated by the increase of the effective size $R$ (defined as per Eq. (9)) as $\gamma$ is approaching $\gamma_{\mathrm{cr}}$.
@@ III.2 Excited-state (ES) solitons
Stable higher-order (ES) soliton solutions have been constructed by applying the imaginary-time-simulation method to Eqs. (2) and (3) with the input taken as per Eq. (15), up to $S_{+}$ (still higher values of $S_{+}$ were not considered here, cf. Refs. [50, 51]). Typical examples of the so density and phase structure of the so constructed stable ES solitons with $S_{+}=1,2,3$ and $4$ are displayed in Fig. 3, where we set $g_{\pm}=\gamma=1$ and $N=4$.
@@
As said above, the stability of the ES-SV solitons was verified through systematic real-time simulations of their perturbed evolution, using the split-step fast-Fourier-transform algorithm. Typical examples of the stable evolution of the solitons with $S_{+}=1,2,3$, and $4$ are displayed in Figs. 5(a1-a4), respectively. In these plots, parameter values $(N,\gamma)$ are chosen from the stability region delineated in Fig. 4. On the other hand, when parameters $(N,\gamma)$ are taken in instability areas of Fig. 4, the simulations demonstrate spontaneous destruction of the SV solitons in the course of the long-time propagation in Figs. 5(b1-b4).
Figure 5: The stable simulated long-real-time evolution of stable ES-SV is displayed by means of density patterns of $\left|\Psi_{+}\right|^{2}$, for $(S_{+},\gamma)=(1,0.6)$ (a1), $(2,0.6)$ (a2), $(3,0.6)$ (a3) and $(4,0.6)$ (a4). Examples of the unstable evolution are presented in panels (b1-b4) for $(S_{+},\gamma)=(1,1.1)$ (b1), $(2,1.4)$ (b2), $(3,1.5)$ (b3) and $(4,1.7)$ (b4). Here, fixed parameters are $g_{\pm}=1$ and $N=6$.
@@
To explain the counter-intuitive phenomenon observed in Fig. 4, where the stability area of the ES-SV solitons expands with then increase of the extra topological charge $S_{+}$, the corresponding ratio $F=N_{-}/N_{+}$ (see Eq. (9)) is displayed in Fig. 6(a) vs. $S_{+}$ for fixed parameters $(N,g_{\pm},,\gamma)=(4,1,1)$. The growing dependence $F\left(S_{+}\right)$ implies the transfer of the condensate into the self-repulsive component $\Psi_{-}$, which naturally inhibits the onset of the collapse.
@@
To further quantify properties of the ES-SV solitons, we fix parameters $(N,g_{-},\gamma)=(4,1,1)$ and vary strength $g_{+}$ of the self-attraction component in component $\Psi_{-}$, plotting the respective dependences $\mu\left(g_{+}\right)$ for different values of $S_{+}$ in Fig. 6(b). An essential observation is that $\mu(g_{+})$ curves satisfy the Vakhitov-Kolokolov (VK) criterion, $d\mu/dg_{+}<0$, which is a well-known necessary stability condition for solitons maintained by a self-attractive nonlinearity [48, 24] (the VK criterion takes this form for the fixed norm, while the strength of the self-attraction varies).
@@
Figure 6: (a) Ratio $F=N_{-}/N_{+}$ (see Eq. (9)) for the ES-SV solitons vs. the additiona; topological charge $S_{+}$, with fixed parameters $(N,g_{+},\gamma)=(4,1,1)$. (b) The solitons’ chemical potential $\mu$, for different values of $S_{+}$, as a function of the self-interaction strength $g_{+}$ in component $\Psi_{+}$, with fixed parameters $(N,g_{-},\gamma)=(4,1,1)$.
@@
Next, Fig. 7 displays the dependence of $\mu$ and effective soliton’s radius $R$ (see Eq. (9)) as functions of the self-repulsion strength $g_{-}$ in component $\Psi_{-}$ for different values of $S_{+}$, fixing other parameters as $(N,g_{+},\gamma)=(4,1,1)$. In particular, positive slope, $d\mu/dg_{-}>0$ of the $\mu\left(g_{-}\right)$ curves in Fig. 7(a) implies that the ES-SV families satisfy the anti-VK criterion, which is a necessary stability condition for solitons with respect to the variation of the self-repulsion strength [49]. In Fig. 7(b), the growth of the soliton’s radius $R(g_{-})$ with the increasing of $g_{-}$ demonstrates a natural trend to expansion of the bound state under the action of stronger self-repulsion.
@@
Figure 7: (a,b) Chemical potential $\mu$ and effective radius $R$ of the ES-SV solitons vs. the self-repulsion strength $g_{-}$ in component $\Psi_{-}$ for different values of the additional topological charge $S_{+}$, with fixed parameters $(N,g_{+},\gamma)=(4,1,1)$.
@@ IV Stability of moving 2D solitons
Typical examples of the stable moving GS-SV solitons with velocity vectors $\left(V_{x},V_{y}\right)=\left(0,0.09\right),$ $\left(0.09,0\right)$, and $\left(0.0636.0.0636\right)$ are presented, severally, in the top, middle, and bottom rows of Fig. 8, for fixed parameters $(N,\gamma,g_{\pm},S_{+})=(4,1,1,0)$.  In this case, $V_{\mathrm{cr}}$ is very close to the velocity value $0.09$ for which the solutions are produced in the top and middle rows of Fig. 8. The velocity corresponding to the bottom row, $V=\sqrt{2}\times 0.06\approx 0.0\allowbreak 85$, is also close to this value.
These results are summarized in Fig. 9 by dependences of $V_{\mathrm{cr}}$ on parameters $g_{+}=g_{-}\equiv g$ and $N$. In particular, Fig. 9(a) shows that, in the range of $3\leq N\leq 5.5$, the critical velocity steeply increases with the growth of $N$. This dependence is explained by the fact that stronger nonlinearity, which corresponds to larger norm, makes the destabilizing effect of the velocity in Eqs. (16) and (17) effectively weaker. The same argument explains the growth of $V_{\mathrm{cr}}$ with the increase of $g$ in Fig. 9(b). The availability of the stably moving SV solitons suggests considering collisions between them. In particular, we have performed simulations for the collisions between ones moving with opposite velocities $V_{x}=\pm 0.03$ in Fig. 10. Figs. 10(a)-10(c) display the contour plots of $\left|\Psi_{+}\right|^{2}+\left|\Psi_{-}\right|^{2}$ at $t=510$, $t=1035$ and $t=1500$, respectively. The arrows in the figures represent the direction of solitons motion. Isosurface plot is displayed in Fig. 10(d). Clearly, the solitons repel each other and maintain their shapes after the rebound, i.e., they collide elastically.
@@
Figure 8: Isosurface and contour plots of the component densities $\left|\Psi_{+}\right|^{2}$ and  $\left|\Psi_{-}\right|^{2}$ (panels (a1), (b1), (c1); (a3), (b3), (c3), and (a2), (b2), (c2); (a4), (b4), (c4), respectively) for stable moving GS-SV solitons, with parameters $(N,g_{\pm},\gamma,S_{+})=(4,1,1,0)$. Panels (a1)-(a4), (b1)-(b4), and (c1)-(c4) represent the solitons with velocities $\left(V_{x},V_{y}\right)=\left(0,0.09\right)$, $\left(V_{x},V_{y}\right)=\left(0.09,0\right)$, and $\left(V_{x},V_{y}\right)=\left(0.0636,0.0636\right),$ respectively.
@@
Figure 10: Collisions between SV solitons moving with opposite velocities, $V_{x}=\pm 0.03$. Contour and isosurface plots of $\left|\Psi_{+}\right|^{2}+\left|\Psi_{-}\right|^{2}$ are displayed at $t=510$ (a), $t=1035$ (b), $t=1500$ (c), and (d) (isosurface plot). Here, parameters of the SV solitons are selected as $(N,g_{\pm},\gamma,S_{+})=(4,1,1,0)$.
Figure 11: An example of a stable mixed mode with parameters $(N,g_{\pm},\gamma,V_{x},V_{y})=(4,1,1,0,0.1)$. (a) One-dimensional cross sections of the two components, $\left|\Psi_{+}\right|$ and $\left|\Psi_{-}\right|$. Panel (b) displays the density patterns of vortex component, $\left|\Psi_{+}\right|^{2}$.
@@ V Conclusion
A natural direction for further research is to extend the current analysis for ESs of solitons of the MM (mixed-mode) type, which combine terms with vorticities 00 and ±$\pm 1$ in the two components. It may be also interesting to explore Josephson oscillations between ESs of solitons of the SV and MM types, cf. Ref. [54]. A challenging possibility is to seek for (meta-) stable ESs of solitons in three-dimensional SOC BEC, cf. Ref. [40].
Recently, much interest in the experimental [55, 56, 13, 14, 15] and theoretical [57, 29, 58, 59, 60, 61] work was drawn to self-bound quantum droplets in BEC stabilized by effects of quantum fluctuations. In this connection, it will be relevant to construct stable ESs of spin-orbit-coupled quantum droplets of the SV and MM types. To this end, it is necessary to use SOC GPEs including the Lee-Huang-Yang corrections [59].

### v2

@@ I Introduction
Various methods have been proposed to achieve this purpose. On the one hand, it has been predicted that long-range interactions in BEC of dipole atoms, carrying permanent magnetic moments, can support stable 2D or even 3D matter-wave solitons [30, 31, 32, 33, 34]. Another possibility is to use the combination of the spin-orbit coupling (SOC) between two components of binary BEC and attractive mean-field nonlinearity [35, 37, 38, 39, 40, 41, 42, 43]. In this connection, it is relevant to mention that SOC has been experimentally realized in a 2D setup [44]. However, the stabilization provided by these systems is limited to the fundamental (ground state, GS) solitons. In particular, the SOC system makes it possible to find 2D excited-state (ES) solitons, which, however, are completely unstable [35]. Naturally, ES solitons exhibit a richer amplitude and phase structure than their GS counterparts, which makes stabilization of ES solitons another relevant problem, that has not been resolved in previous works [26, 45].
@@
In the experiment, SVs (alias half-vortices [36]) can be induced by vorticity-carrying laser beams resonantly coupled to particular components of the binary BEC. Our analysis demonstrates that the ES solitons may be stable, at least, up to $S_{+}=6$. Systematically collected numerical results reveal that the ES solitons with the vorticity in both components exist provided that their total does not exceed a collapse threshold $N_{\mathrm{thr}}$, which surpasses the well-known threshold ($N_{c}\approx 5.85$) for the fundamental (Townes) single-component solitons in the free 2D space [21, 22]. The value $N_{\mathrm{thr}}$ increases with the growth of the additional vorticity $S_{+}$. These findings offer, for the first time, a solution of the problem of the creation of stable ES solitons in SOC systems.
The following presentation is structured as follows. In Section II, we introduce the model, based on a system of two spin-orbit-coupled Gross-Pitaevskii equations (GPEs), with the opposite signs of the cubic nonlinearity in the two components. In Section III, we produce SV soliton solutions of the GS and ES types and test their stability by means of numerical methods. In Section IV, we address moving SV solitons and their stability, which is a nontrivial problem, as the SOC breaks the system’s Galilean invariance. A critical velocity $V_{\mathrm{cr}}$, up to which the stable moving solitons exist, is identified as a function of the total norm $N$. Finally, Section V concludes the paper.
@@ II The model
where positive $g_{+}$ and $g_{-}$ coefficients are strengths of the self-attraction and repulsion in the components $\Psi_{+}$ and $\Psi_{-}$, respectively, and $\gamma$ is the cross-interaction strength. Note that the form in which the equations are written, with the fixed value of the SOC coefficient, does not admit rescaling of the coordinates. Rescaling of the wave function is possible, which may be used to fix that any of coefficients $g_{\pm}$ and $\gamma$ is $1$. In fact, in most cases, we use this degree of freedom to set $\gamma=1$, except for the solutions presented in Figs. 2(a,b,c), where $\gamma$ is varied, while the nomalization is fixed by setting $g_{\pm}=1$.
The linearized form of Eqs. (2) and (3) for plane-wave solutions, $\Psi_{\pm}\sim\exp\left(-i\mu_{\mathrm{linear}}t+ik_{x}x+ik_{y}y\right)$, yields the dispersion relation
@@
Obviously, $N$ is a dynamical invariant of the system of Eqs. (2) and (3). In all cases, when the scaling is used to set either $\gamma=1$ or $g_{\pm}=1$, as said above, $N$ remains a free control parameter of solution families. In the next section, we first set $g_{+}=g_{-}=1$, while the case of $g_{+}\neq g_{-}$ is considered at the section.
@@
is also a dynamical invariant. Note that the  substitution of ansatz (6) in expression (13) yields a simple relation between the angular momentum and norm (10) of the stationary states, which applies equally well to the GS and ESs:
@@ III.1 Ground-state (GS) solitons
For two typical values of the total norm, $N=4.5$ and $N=3.5$, properties of the GS solutions are summarized in Figs. 2(a,b,c), where the chemical potential, effective radius, and the norm ratio of the two components, defined as per Eq. (9), are plotted vs. the cross-interaction strength $\gamma$ in Eqs. (2) and (3). In these plots, $\gamma$ takes both positive and negative values, which correspond, respectively, to the cross-attraction and repulsion, while the self-interaction coefficients are fixed as $g_{+}=g_{-}=1$. In Fig. 2(a), the SV soliton suffers delocalization, turning into an infinitely broad state with an infinitely small amplitude, as chemical potential $\mu$ is reaching the limit value (5), $\mu_{\mathrm{cutoff}}=-1/2$, at $\gamma\rightarrow\gamma_{\mathrm{cr}}\left(N=4.5\right)=-0.775$ or $\gamma_{\mathrm{cr}}\left(N=3.5\right)=-0.24$. For other values of the norm (in the interval of $5.53<N<$ 5.5), the dependence of $\gamma_{\mathrm{cr}}$ on $N$ is plotted in Fig. 2(d). Self-trapped bound states (solitons) do not exist if $\gamma$ is “too negative”, taking values $\gamma<\gamma_{\mathrm{cr}}(N)$, as the self-trapping is obviously impossible when the overall nonlinearity is self-repulsive. The transition to the delocalization is also demonstrated by the increase of the effective size $R$ (defined as per Eq. (9)) as $\gamma$ is approaching $\gamma_{\mathrm{cr}}$.
@@ III.2 Excited-state (ES) solitons
Stable higher-order (ES) soliton solutions have been constructed by applying the imaginary-time-simulation method to Eqs. (2) and (3) with the input taken as per Eq. (15), up to $S_{+}=6$ (still higher values of $S_{+}$ were not considered here, cf. Refs. [52, 53]). Typical examples of the so density and phase structure of the so constructed stable ES solitons with $S_{+}=1,2,3$ and $4$ are displayed in Fig. 3, where we set $g_{\pm}=\gamma=1$ and $N=4$.
@@
To explain the counter-intuitive phenomenon observed in Fig. 4, where the stability area of the ES-SV solitons expands with then increase of the extra topological charge $S_{+}$, the corresponding ratio $F=N_{-}/N_{+}$ (see Eq. (9)) is displayed in Fig. 5(a) vs. $S_{+}$ for fixed parameters $(N,\gamma,g_{\pm})=(4,1,1)$. The growing dependence $F\left(S_{+}\right)$ implies the transfer of the condensate into the self-repulsive component $\Psi_{-}$, which naturally inhibits the onset of the collapse.
@@
To further quantify properties of the ES-SV solitons, we fix parameters $(N,\gamma,g_{-})=(4,1,1)$ and vary strength $g_{+}$ of the self-attraction component in component $\Psi_{-}$, plotting the respective dependences $\mu\left(g_{+}\right)$ for different values of $S_{+}$ in Fig. 5(b). An essential observation is that $\mu(g_{+})$ curves satisfy the Vakhitov-Kolokolov (VK) criterion, $d\mu/dg_{+}<0$, which is a well-known necessary stability condition for solitons maintained by a self-attractive nonlinearity [50, 24] (the VK criterion takes this form for the fixed norm, while the strength of the self-attraction varies).
@@
Figure 5: (a) Ratio $F=N_{-}/N_{+}$ (see Eq. (9)) for the ES-SV solitons vs. the additional topological charge $S_{+}$, with fixed parameters $(N,\gamma,g_{+})=(4,1,1)$. (b) The solitons’ chemical potential $\mu$, for different values of $S_{+}$, as a function of the self-interaction strength $g_{+}$ in component $\Psi_{+}$, with fixed parameters $(N,\gamma,g_{-})=(4,1,1)$.
@@
Next, Fig. 6 displays the dependence of $\mu$ and effective soliton’s radius $R$ (see Eq. (9)) as functions of the self-repulsion strength $g_{-}$ in component $\Psi_{-}$ for different values of $S_{+}$, fixing other parameters as $(N,\gamma,g_{+})=(4,1,1)$. In particular, positive slope, $d\mu/dg_{-}>0$ of the $\mu\left(g_{-}\right)$ curves in Fig. 6(a) implies that the ES-SV families satisfy the anti-VK criterion, which is a necessary stability condition for solitons with respect to the variation of the self-repulsion strength [51]. In Fig. 6(b), the growth of the soliton’s radius $R(g_{-})$ with the increasing of $g_{-}$ demonstrates a natural trend to expansion of the bound state under the action of stronger self-repulsion.
@@
Figure 6: (a,b) Chemical potential $\mu$ and effective radius $R$ of the ES-SV solitons vs. the self-repulsion strength $g_{-}$ in component $\Psi_{-}$ for different values of the additional topological charge $S_{+}$, with fixed parameters $(N,\gamma,g_{+})=(4,1,1)$.
@@ IV Stability of moving 2D solitons
The dependences of $V_{\mathrm{cr}}$ on parameters $g_{+}=g_{-}\equiv g$ and $N$ are summarized in Fig. 7. In particular, Fig. 7(a) shows that, in the range of $3\leq N\leq 5.5$, the critical velocity steeply increases with the growth of $N$. This dependence is explained by the fact that stronger nonlinearity, which corresponds to larger norm, makes the destabilizing effect of the velocity in Eqs. (16) and (17) effectively weaker. The same argument explains the growth of $V_{\mathrm{cr}}$ with the increase of $g$ in Fig. 7(b). The availability of the stably moving SV solitons suggests considering collisions between them. In particular, we have performed simulations for the collisions between ones moving with opposite velocities $V_{x}=\pm 0.03$ in Fig. 8. Figs. 8(a)-8(c) display the contour plots of $\left|\Psi_{+}\right|^{2}+\left|\Psi_{-}\right|^{2}$ at $t=510$, $t=1035$ and $t=1500$, respectively. The arrows in the figures represent the direction of solitons motion. Isosurface plot is displayed in Fig. 8(d). Clearly, the solitons repel each other and maintain their shapes after the rebound, i.e., they collide elastically.
@@
Figure 8: Collisions between SV solitons moving with opposite velocities, $V_{x}=\pm 0.03$. Contour and isosurface plots of $\left|\Psi_{+}\right|^{2}+\left|\Psi_{-}\right|^{2}$ are displayed at $t=510$ (a), $t=1035$ (b), $t=1500$ (c), and (d) (isosurface plot). Here, parameters of the SV solitons are selected as $(N,\gamma,g_{\pm},S_{+})=(4,1,1,0)$.
Figure 9: An example of a stable mixed mode with parameters $(N,\gamma,g_{\pm},V_{x},V_{y})=(4,1,1,0,0.1)$. (a) One-dimensional cross sections of the two components, $\left|\Psi_{+}\right|$ and $\left|\Psi_{-}\right|$. Panel (b) displays the density patterns of vortex component, $\left|\Psi_{+}\right|^{2}$.
@@ V Conclusion
A natural direction for further research is to extend the current analysis for ESs of solitons of the MM (mixed-mode) type, which combine terms with vorticities 00 and ±$\pm 1$ in the two components. It may be also interesting to explore Josephson oscillations between ESs of solitons of the SV and MM types, cf. Ref. [56]. A challenging possibility is to seek for (meta-) stable ESs of solitons in three-dimensional SOC BEC, cf. Ref. [41] (it is relevant to expect metastability in that case, as the supercritical collapse cannot be suppressed in 3D, hence 3D solitons cannot realize a ground state). Another potentially interesting direction is the study of the stable ESs of solitons in a heteronuclear mixture [57], if SOC is induced in one or both species which compose the mixture.
The underlying system of equations (2) and (3) does not include a trapping potential. It is well known that 2D potentials stabilize all trapped single-component solitons with zero vorticity, $S=0$, against the critical collapse, and also stabilize a part of solitons with embedded vorticity $S=1$ against the splitting. On the other hand, all the trapped solitons with $S\geq 2$, which may be considered as counterparts of the ESs considered here, remain unstable [26]. In terms of the present consideration, the inclusion of a trapping potential in Eqs. (2) and (3) may help to additionally stabilize the ESs.
It is relevant to mention that the stabilization mechanism provided by SOC should be sufficiently stable against the action of losses, i.e., gradual decay of the solitons’ amplitude, as the SOC is, by itself, the linear phenomenon. The topological charge of SV states provides their protection against other external perturbations.
Recently, much interest in the experimental [58, 59, 13, 14, 15] and theoretical [27, 28, 60, 29, 61, 62, 63, 64] work was drawn to self-bound quantum droplets in BEC stabilized by quantum fluctuations (the Lee-Huang-Yang effect [65]). In this connection, it will be relevant to construct stable ESs of spin-orbit-coupled quantum droplets of the SV and MM types. To this end, it is necessary to use SOC GPEs including the Lee-Huang-Yang corrections [62].